The sum of the digits in a 2 digit number is 13. The tens are greater than the ones so what is the digit number

Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20060)

94, 85, and 76 all three have sum of digits 13 and have the tens digit
greater than the ones digit.

Answer by stanbon(75887)

The sum of the digits in a 2 digit number is 13. The tens are greater than the ones so what is the digit number
----------------------
Equations:
Let the number be 10t+u where t is the ten's digit and u is the unit's dig.
----
t + u = 13
t - u > 0
----
2t > 13
===
t 7 pr 8 pr 9
---
If t = 7 then 10t+u = 76
If t = 8 then 10t+u = 85
If t = 9 then 10t+u = 94
===============================
Cheers,
Stan H.